OpenStudy (anonymous):

Write each equation in standard form using integers. y=3x-11 y=-2x+5 y= 5/2x+4 y= 2/3x-25/3

OpenStudy (anonymous):

Standard form is Ax+By=C, where A is non-negative and there are no fractions. This basically means you have to get the x and y terms on the same side, multiply everything by -1 if A is negative, and eliminate fractions. I'll walk you through the third one. y = 5/2x + 4 We want the x and y terms on the same side, with the constant on the opposite. -5/2x + y = 5/2x-5/2x + 4 -5/2x + y = 4 Remember that general form is Ax+Bx=C. A, here, equals -5/2. There are two problems: There is a fraction in the equation, and A is negative. We can kill two birds with one stone by multiplying everything by the LCD (least common denominator, in the case 2) times -1. LCD*-1=-2 here. -2(-5/2x + y) = -2(4) 10/2x + -2y = -8 5x - 2y = -8 That's your answer for #3. I'll leave the rest up to you (or other people).